OpenStudy (anonymous):

what percent of 50000 is 735

OpenStudy (neer2890):

x% of 50000=735 \[\frac{ x*50000 }{ 100 }=735\] find x

OpenStudy (anonymous):

i am cunfused i will give you a meadle if you will tell me the answer plaese

OpenStudy (anonymous):

@neer2890

OpenStudy (neer2890):

x=735/500

OpenStudy (neer2890):

1.47%